OpenStudy (anonymous):

identify the solutions of 3x2 − 13x + 4 = 0

jimthompson5910 (jim_thompson5910):

Do you know the quadratic formula?

hero (hero):

Is the quadratic formula really needed here?

hero (hero):

Why can't we just try to factor it?

jimthompson5910 (jim_thompson5910):

Factoring works, but I like the quadratic formula since that works with any quadratic

jimthompson5910 (jim_thompson5910):

and you don't need to guess and check

hero (hero):

Factoring works too. I see what you mean about guess and check. I don't like it either.

hero (hero):

Completing the square is another way to avoid guess and check

OpenStudy (anonymous):

okay... so what do i do? lol

hero (hero):

I've always viewed the quadratic formula as something you use when all other methods have failed.

hero (hero):

Hint: Split the middle term to get: 3x^2 - 12x - 1x + 4 = 0 Then factor by grouping.

OpenStudy (anonymous):

3x-1 and x-4?

jimthompson5910 (jim_thompson5910):

that's correct, so 3x^2 − 13x + 4 = 0 factors to (3x-1)(x-4) = 0 What's next?

OpenStudy (anonymous):

umm.... i dont knoww

jimthompson5910 (jim_thompson5910):

If AB = 0 then either A = 0 or B = 0

jimthompson5910 (jim_thompson5910):

So if (3x-1)(x-4) = 0 then 3x-1 = 0 or x-4 = 0

OpenStudy (anonymous):

okayyy?

jimthompson5910 (jim_thompson5910):

do you see how I got that?

jimthompson5910 (jim_thompson5910):

Basically, if two numbers multiply to zero, then one of them has to be zero

OpenStudy (anonymous):

yea.. but none of the answer choices is zero

jimthompson5910 (jim_thompson5910):

that's not the final answer

jimthompson5910 (jim_thompson5910):

What do you get when you solve 3x-1 = 0?

OpenStudy (anonymous):

1 over 3

jimthompson5910 (jim_thompson5910):

good, what about x-4 = 0?

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

so the solutions are x = 1/3 or x = 4

OpenStudy (anonymous):

so the final answer is x=1/3 and x=4?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thankss

jimthompson5910 (jim_thompson5910):

np